Circular No. 6204

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SATURN S. J. O’Meara, Sky & Telescope, reports that visual observations by

W. Sheehan, D. Graham, T. Dobbins, and himself with the Lick Observatory 0.91-m refractor show two white spots in the northern region of the equatorial zone. The larger, low-contrast spot of diameter ∼ 4′′ transited the planet at Aug. 10.375 UT (corresponding to system-I longitude 333^{o}). The smaller spot of diameter ∼ 2′′ transited at Aug. 10.444 (longitude 31^{o}). Observations made ∼ 12 hr prior to crossing the ring plane on Aug. 10.5 showed the rings still visible to a distance from the planet of one Saturn diameter.

COMET 19P/BORRELLY P. Lamy, Laboratoire d’Astronomie Spatiale, Marseille, and his team

report: “Using the Wide Field Planetary Camera 2 of the Hubble Space Telescope, we have detected a highly elongated nucleus rotating with a synodic period of 24.7 hr. The prolate spheroid that gives the best fit to the nuclear-magnitude light curve has major and minor axes dimensions of 8.3 and 3.3 km, respectively, assuming a geometric albedo of 4 percent. We estimate that about 10 percent of the surface area is active.”

GRS 1915+105 B. A. Harmon, W. S. Paciesas, and G. J. Fishman, Marshall Space

Flight Center, NASA, report, for the Compton Observatory BATSE Team: “The 20- to 100-keV hard-x-ray flux from GRS 1915+105, measured by BATSE, has been increasing gradually since late July. The source showed variability on a 1-day timescale; the average flux was ∼ 150 mCrab for Aug. 5–8. This sustained level of activity is comparable to that seen in 1995 Jan.–Feb.”

F. D. Ghigo, National Radio Astronomy Observatory; and E. B. Waltman and R. S. Foster, Naval Research Laboratory (NRL), write: “The NRL-Green Bank Interferometer Monitoring Program has detected a radio flare in the galactic superluminal jet source 1915+105. Flux densities reached 750 mJy at 2 GHz and 400 mJy at 8 GHz at Aug. 10.04 UT, remained at this level through Aug. 10.21, and declined to 400 mJy (2 GHz) and 125 mJy (8 GHz) at Aug. 10.98.”

1995 August 11 Daniel W. E. Green